OpenStudy (anonymous):

Solve. 8^2-4(-8)(-7)

jimthompson5910 (jim_thompson5910):

what is 8^2 ?

OpenStudy (anonymous):

16

jimthompson5910 (jim_thompson5910):

8^2 = 8*8 = ???